# 1bit-Merging: Dynamic Quantized Merging for Large Language Models

### v1

@@ 4.1 Experimental Setup
Benchmark.
Our experimental evaluation focuses on three model families: LLaMA-2 7B Touvron et al. (2023), Mistral 7B Jiang et al. (2023), and LLaMA-2 13B Touvron et al. (2023), each covering distinct specializations in: general knowledge (Chat), mathematical reasoning (Math), and code generation (Code). We assess performance using seven benchmark datasets across three domains: MMLU Hendrycks et al. (2020), HellaSwag Zellers et al. (2019) and TruthfulQA Lin et al. (2022) for assessing general knowledge and reasoning capabilities; GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021) for testing mathematical reasoning proficiency; and HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) for evaluating code generation abilities. To ensure consistent and unbiased assessment, model performance is evaluated using zero-shot accuracy, with pass@1 rate specifically measuring code generation correctness.
4.2 Main Results
Using the Chat111huggingface.co/meta-llama/Llama-2-7b-chat-hf, Math222huggingface.co/TIGER-Lab/MAmmoTH-7B, and Code333huggingface.co/mrm8488/llama-2-coder-7b fine-tuned models derived from the same LLaMA2-7B pretrained model444huggingface.co/meta-llama/Llama-2-7b-hf, we first compare the performance of the compressed expert models with the individually fine-tuned models to demonstrate the effectiveness of our task vector compression. Subsequently, we evaluate the performance of our 1bit-Merging model against existing task-vector-based model merging methods. Table 3 presents a comprehensive comparison across seven datasets.
@@ 4.2 Main Results
Compressed Expert Models.
Our task vector compression method demonstrates remarkable effectiveness, with the compressed expert models not only maintaining but often surpassing the performance of their fine-tuned counterparts. The 1bit-Chat model achieves an average score of 25.55, surpassing its fine-tuned counterpart’s 25.50. More notably, the 1bit-Math model improves from 31.52 to 32.17, while the 1bit-Code model shows the most substantial gain, increasing from 24.98 to 25.97. Moreover, the compressed expert models exhibit strong adaptability, maintaining high performance not only in their specialized domains but also across other non-specialized domains. (1) Superior performance in specialized domains: Within their specialized domains, the 1bit-Chat model enhances general knowledge capabilities with a 0.63 point improvement, the 1bit-Math model maintains near-optimal performance, and most impressively, the 1bit-Code model achieves a substantial 3.0 point boost in code generation proficiency. These results suggest that our compression method not only preserves but often enhances domain-specific expertise. (2) Superior performance in non-specialized domains: The benefits of our compression technique extend beyond primary specializations, showcasing enhanced cross-domain generalization. This is particularly evident in the compressed Math model, which achieves notable gains of 1.11 points in general knowledge and 1.25 points in code generation—domains outside its primary expertise. These consistent improvements across both specialized and general domains indicate that our compression method facilitates positive knowledge transfer.
Merging Models with 1bit-Merging.
Our 1bit-Merging method then integrates the compressed expert models through dynamic routing. As shown in Table 3, our 1bit-Merging method demonstrates superior performance compared to existing merging techniques in two significant aspects. (1) Consistent improvements across all tasks: While baseline methods exhibit distinctive domain-specific strengths - with Ties-Merging excelling in mathematical reasoning and DARE showing superior results in general knowledge and code generation - our 1bit-Merging method achieves comparable or better performance across all domains. Specifically, our 1bit-Merging method surpasses Ties-Merging by 1.45 points in mathematical reasoning and DARE by 2.71 points in general knowledge. Although we observe a modest 0.75-point decrease in code generation compared to DARE, this minor regression is primarily attributable to the lower baseline performance of the Code fine-tuned model. Notably, our method still achieves a substantial 8.35-point improvement over the original code fine-tuned model. (2) Exceeds the highest average performance of expert fine-tuned models: Perhaps most significantly, our method achieves a remarkable average performance score of 36.74 across all tasks, markedly surpassing the average best performance (35.16) of task-specific fine-tuned models—a distinctive achievement that none of the existing merging methods have accomplished. This remarkable improvement of 1.58 points over the highest individual expert scores demonstrates that our 1bit-Merging method not only preserves but actually enhances the specialized capabilities of individual models, suggesting the merged model leverages complementary strengths across different domains and maintains their peak performance.
General Knowledge Mathmetical Reasoning Code Generation
/ 59.05 65.97 55.69 42.53 9.16 49.6 42.7 46.37
Attention 53.25 57.03 54.59 39.20 8.14 46.6 36.0 42.12
MLP 54.71 58.00 53.24 38.51 8.14 47.6 35.4 42.23
Chat Linear 54.95 58.83 53.73 38.82 8.26 47.6 37.2 42.77
/ 60.77 58.68 44.68 63.38 22.74 38.1 23.8 44.59
Attention 60.11 58.48 41.13 58.76 19.92 41.1 25.6 43.59
MLP 58.74 58.12 43.57 53.37 16.68 44.9 29.3 43.53
Math Linear 57.93 58.24 39.41 47.16 13.9 41.1 25.6 40.48
/ 50.58 53.19 45.29 31.69 4.84 50.9 40.9 39.63
Attention 47.88 53.51 42.96 27.29 5.16 41.6 33.5 35.99
MLP 45.65 54.56 42.59 34.34 6.08 47.9 28.0 37.02
Code Linear 39.99 53.80 36.47 22.52 4.82 40.6 31.1 32.76
Table 5: Impact of different layer types on compression performance on Mistral 7B models.
4.3 Using Different Model Architecture
To validate the generalizability of our method, we extend our experiments to the Mistral 7B architecture. As illustrated in Table 3, Our approach demonstrates robust performance improvements despite architectural differences. When merging Chat555huggingface.co/mistralai/Mistral-7B-Instruct-v0.1, Math666huggingface.co/TIGER-Lab/MAmmoTH2-7B, and Code777huggingface.co/Nondzu/Mistral-7B-codealpaca-lora fine-tuned models derived from pretrained backbone 888huggingface.co/mistralai/Mistral-7B-v0.1. 1bit-Merging outperforms Task Arithmetic by an average of 9.26 points, Ties-Merging by 2.13 points, and DARE by 3.46 points across all evaluated datasets. These enhancements are particularly significant in mathematical reasoning and general knowledge tasks. On the benchmark datasets MATH and GSM8K, our method respectively achieves accuracies of 20.60% and 60.42%, significantly exceeding DARE—the strongest baseline for mathematical reasoning—by 5.60 points (a 33.73% relative improvement) on MATH and by 3.03 points (a 5.28% relative improvement) on GSM8K. Additionally, on the TruthfulQA dataset, 1bit-Merging demonstrates substantial gains over all baselines (12.06 points over Task Arithmetic, 4.35 points over Ties-Merging, and 10.22 points over DARE) while performing on par with the Chat fine-tuned model. This underscores the effectiveness of 1bit-Merging in maintaining the peak performance in fine-tuned expert models. Furthermore, the average performance of our 1bit-Merging method across all datasets is 48.43, surpassing the strongest Chat fine-tuned model’s average of 46.37—a distinction not achieved by any other baseline methods.
@@ 4.3 Using Different Model Architecture
Table 6: Impact of different layer types on compression performance on LLaMA-2 13B models.
4.4 Scaling to Larger Model Size
We further evaluate the scalability of our method using LLaMA-2 13B999huggingface.co/meta-llama/Llama-2-13b-hf architecture by merging Chat101010huggingface.co/meta-llama/Llama-2-13b-chat-hf, Math111111huggingface.co/TIGER-Lab/MAmmoTH-13B, and Code121212huggingface.co/emre/llama-2-13b-code-chat fine-tuned models. As shown in Table 4, our approach remains effective at this larger scale. The compressed expert models demonstrate consistent enhancement, achieving an average improvement of 0.54 points over their original fine-tuned counterparts. Specifically, the compressed experts show consistent improvements across all specialized domains: 1bit-Chat achieves a 0.11-point gain in general knowledge tasks, 1bit-Math demonstrates a 0.38-point improvement in mathematical reasoning, and 1bit-Code shows a notable 1.05-point enhancement in code generation. Additionally, our 1bit-Merging method demonstrates superior performance against all baseline approaches, surpassing Task Arithmetic by 6.12 points, Ties-Merging by 1.18 points, and DARE by 0.89 points on average across all evaluated datasets. When compared to DARE, the strongest baseline for mathematical reasoning and general knowledge, our method achieves notable improvements of 1.32 points (14.54% relative increase) on MATH and 1.70 points (4.20% relative increase) on TruthfulQA. In code generation tasks, our method outperforms the leading baseline, Ties-Merging, by an average of 0.45 points.
Table 7: Performance comparison between merging original and compressed fine-tuned LLaMA2-7B models.
MMLU HellaSwag TruthfulQA GSM8K MATH MBPP HumanEval
LLaMA2-7B 41.50 49.63 37.45 47.34 6.46 13.5 7.3 29.03
Mistral-7B 47.34 46.80 41.00 52.16 13.26 32.1 29.9 39.17
LLaMA2-13B 52.22 57.52 41.49 49.89 7.32 24.1 9.1 34.95
4.5 Ablation Studies
Layer Types on Compression Performance.
We further evaluate how compression position affects the performance of compressed expert models across Mistral 7B and LLaMA-2 13B architectures. As shown in Tables 5 and 6, for Chat models, while the optimal quantization locations showed slight variations—with Mistral 7B preferring all Linear layers and LLaMA2 13B favoring attention layers—the performance differences between these choices were minimal. However, for specialized Math and Code models, we find quantizing attention layers consistently to be most effective in preserving domain-specific capabilities.
Merging Compressed Expert Models.
Merging compressed expert models proves effective for Task Arithmetic but shows limitations with parameter-dropping methods like Ties-Merging and DARE. This is because parameter-dropping techniques destroy the compressed knowledge distributions essential for specialized task performance. In our experiments with LLaMA2-7B and LLaMA2-13B, where compressed experts show marginal improvements over individually fine-tuned models, merging compressed experts yields performance gains of 1.06 and 1.14 points respectively, compared to merging their fine-tuned counterparts.
Figure 3: Performance and storage trade-off of different methods in merging fine-tuned Mistral 7B models.
4.6 Performance vs. Storage Trade-offs
We demonstrate the trade-off between performance and storage requirements in Figure 3. We use Mistral 7B models, as its Code model demonstrates superior performance in code generation. Task-specific routing (Routing) achieves performance equivalent to individually fine-tuned models, serving as a performance upper bound. However, Routing requires maintaining full parameters for each task model, leading to substantial storage overhead. While model merging methods reduce storage requirements, they underperform Routing by over 4.93 points. Our 1bit-Merging strikes a favorable balance between these extremes, maintaining 94.53% of Routing performance while requiring only 66.25% of its storage. Furthermore, leveraging the observation that Chat models exhibit minimal sensitivity to compression positions, we applied quantization on all Linear layers for the Chat model. This optimization achieves an even more efficient storage reduction to 55.02% while preserving comparable performance levels.
5 Conclusion
Limitations
Despite the promising results of 1bit-Merge, several limitations deserve attention. First, while our method significantly reduces storage requirements, the dynamic routing mechanism introduces computational overhead during inference, as the router must analyze each input to determine the optimal task vector and perform binary transformations. Although this overhead is relatively small, it could impact real-time applications with strict latency requirements.
Second, while compressed expert models show enhanced performance within their specialized domains, our framework may not fully capture complex cross-task interactions that could enable more sophisticated knowledge transfer, particularly for tasks requiring simultaneous expertise from multiple domains.
Third, while our 1-bit quantization approach preserves model performance for most tasks, it may not be optimal for all types of task-specific knowledge, particularly for tasks requiring high numerical precision or fine-grained reasoning.
Ethics Statement

### v2

@@ 4.2 Compare to Expert Models and Traditional Merging
4.3 Compare to Dynamic Merging Method
We conduct additional experiments with Twin-Merging Lu et al. (2024), which utilizes SVD compression for task vectors and, similar to our method, employs dynamic merging to select the appropriate expert model based on input characteristics. To ensure a comprehensive and fair comparison, we set the rank $r=128$, matching the memory storage of 1-bit quantization. This configuration allows for a direct comparison between compression techniques under equivalent storage constraints. We apply both methods to the Chat, Math, and Code expert models derived from LLaMA2-7B, and evaluate their performance after merging.
The performance of compressed expert models using both 1-bit quantization and SVD compression on task vectors is summarized in Table 4. Our results show that SVD compression fails to fully capture fine-tuning information and consistently underperforms 1-bit quantization across all expert models. Notably, low-rank approximation leads to significant degradation when compressing math and code-related task vectors. For instance, SVD-compressed Math models exhibit a 1.97-point drop in GSM8K accuracy, while Code models show a 6.1-point reduction in HumanEval compared to their 1-bit quantized counterparts. As a result, combining SVD-compressed vectors with dynamic routing struggles to match the effectiveness of our 1bit-merging approach, as the fundamental information loss during compression limits the potential benefits of the routing mechanism.
1bit-Merging outperforms Twin-Merging by 2.09 points on GSM8K (48.52 vs. 46.43), as well as the higher MATH score (9.04 vs. 8.14), indicating stronger mathematical reasoning capabilities. In code generation, the performance gap is even more pronounced: 1bit-Merging improves upon Twin-Merging by 5.8 points in HumanEval (18.9 vs. 13.1) and 1.0 points in MBPP (30.1 vs. 29.1). While Twin-Merging shows slightly better general knowledge performance on MMLU (48.36 vs. 47.23), our method maintains competitive results overall and surpasses Twin-Merging in task-specific domains that require precise reasoning.
MMLU HellaSwag TruthfulQA GSM8K MATH MBPP HumanEval
LLaMA2-7B 41.50 49.63 37.45 47.34 6.46 13.5 7.3 29.03
Mistral-7B 47.34 46.80 41.00 52.16 13.26 32.1 29.9 39.17
LLaMA2-13B 52.22 57.52 41.49 49.89 7.32 24.1 9.1 34.95
Table 7: Performance comparison between merging original and compressed fine-tuned models.
4.4 Scalability to Architecture and Size
Generalization to Different Architectures
To assess the architectural generality of our method, we apply 1bit-Merging to expert models based on Mistral 7B. As shown in Table 3 and Table 6, our method consistently performs well across domains. 1bit-Merging outperforms merging strategies—Task Arithmetic (+9.26), TIES-Merging (+2.13), and DARE (+3.46) on average across seven datasets. For mathematical reasoning, 1bit-Merging substantially surpasses TIES-Merging by 4.92 points on GSM8K and 5.6 points on MATH, demonstrating its strong capability in preserving reasoning ability.
Scaling to Larger Model Size
To evaluate the scalability of our method, we apply 1bit-Merging to three fine-tuned variants of LLaMA-2 13B: Chat, Math, and Code. As shown in Table 6 and Table 11, our approach remains effective at this larger scale, outperforming all baseline methods on average—by 6.12 points over Task Arithmetic, 1.18 points over TIES-Merging, and 0.89 points over DARE. Notably, 1bit-Merging achieves a 14.54% relative improvement (1.32 points) over DARE on MATH and a 4.20% relative gain (1.70 points) on TruthfulQA. In code generation, it consistently surpasses TIES-Merging by 0.45 points across MBPP and HumanEval.
5 Further Analyses
Ablation Studies
Merging compressed expert models is effective for Task Arithmetic, but shows limitations when applied to parameter-dropping methods such as TIES-Merging and DARE (see Table 7). This is likely due to the fact that parameter-dropping techniques disrupt the structured knowledge distributions critical for task-specific performance. In our experiments on LLaMA2-7B and LLaMA2-13B, we observe that compressed expert models already achieve slightly better performance than their individually fine-tuned counterparts. Merging these compressed experts further improves performance by 1.06 and 1.14 points, respectively, over merging the fine-tuned models.
Figure 3: Performance vs. storage trade-offs for Mistral 7B deployment. Task-specific routing (Routing) achieves strong performance but requires full parameter storage. Model merging reduces storage at the cost of performance. 1bit-Merging strikes a better balance, and applying 1-bit quantization to all linear layers in Chat (“w/ Chat on Linear”) further improves efficiency.
Performance vs. Storage Trade-offs
Figure 3 illustrates the trade-off between performance and storage requirements among different model deployment strategies, using Mistral 7B expert models. Task-specific routing (Routing), which maintains full parameters for each expert, achieves performance comparable to individually fine-tuned models and serves as an upper bound. However, this comes at a significantly increased storage cost. Model merging methods reduce storage demands, but generally underperform Routing by more than 4.93 points on average. In contrast, our 1bit-Merging strikes a favorable balance—retaining 94.53% of the performance of Routing while requiring only 66.25% of its storage. Notably, we observe that Chat models are less sensitive to the choice of quantized modules. Leveraging this, we apply 1-bit quantization to all linear layers in the Chat model, further reducing storage usage to 55.02% with minimal impact on performance.
6 Conclusion
Limitations
Our approach faces limitations that are both shared with existing merging methods. (1) At a theoretical level, despite the empirical success of weight interpolation techniques, we still lack a comprehensive understanding of why and when these methods work effectively. Although recent research has revealed interesting patterns about weight disentanglement and cross-task linear relationships, the precise conditions for optimal weight interpolation remain unclear. (2) A practical constraint lies in the architectural requirements of our method. Currently, the merging process is restricted to models sharing identical architectures, as it relies on direct parameter operations between models. This limitation becomes particularly challenging when seeking to merge models with specific capabilities, as finding suitable fine-tuned models with matching architectures can be difficult. Looking ahead, extending model merging capabilities to heterogeneous architectures represents an exciting and challenging frontier for future research, though this would require developing novel techniques for mapping and aligning parameters across different architectural designs.
Ethics Statement
Appendix A Appendix
A.1 Implementation Details
Router
We implement the router as a three-layer linear network with Leaky ReLU activations and batch normalization, serving as an input classifier that categorizes each input instance into one of the three task types: math, code, or chat. To train the router, we construct a dedicated validation dataset by randomly sampling instances from the respective downstream datasets and combining them. This ensures that the training data is representative of the diverse input distributions the router is expected to encounter during inference. For the training process, we used a learning rate of 5e-5 and trained the model for 20 epochs on the validation set. The validation set was taken from a split of the overall training set, and we used 800 samples for each downstream dataset during the router training.
Hyperparameters
For Task Arithmetic, we use a default scaling coefficient of $\lambda=1$, which maintains the original magnitude of task vector when adding the pretrained backbone. However, the DARE method has been observed to be more sensitive to variations in both the scaling coefficient $\lambda$ and the drop rate parameter $r$. To achieve a balanced performance, we set the scaling coefficient to $\lambda=0.5$ and establish a default drop rate of $r=0.5$ for DARE. Similarly, for TIES-Merging, which requires the specification of a masking ratio, we set the default mask ratio to $r=0.7$ across all experiments.
A.2 Model Variants and Sources
We evaluate our method across multiple fine-tuned variants of LLaMA2-7B, Mistral 7B, and LLaMA2-13B models, focusing on specialized capabilities in instruction-following (Chat), mathematical reasoning (Math), and code generation (Code). Table 8 summarizes the fine-tuned variants used for LLaMA2-7B along with their corresponding Hugging Face identifiers and use cases. For completeness, we also include similarly fine-tuned versions from other base models: Mistral 7B (Table 9) and LLaMA2-13B Table 10. These variants are obtained from publicly available checkpoints and are used consistently across all compression and merging experiments reported in this work.
Model Type Hugging Face Identifier Use Case
LLaMA2-7B Pretrained Base Model meta-llama/Llama-2-7b-hf General language modeling
Chat-Tuned Model meta-llama/Llama-2-7b-chat-hf Dialogue and instruction-following
Math-Tuned Model TIGER-Lab/MAmmoTH-7B Mathematical reasoning
Code-Tuned Model mrm8488/llama-2-coder-7b Code generation
Table 8: Fine-tuned variants and corresponding Hugging Face identifiers for LLaMA-2 7B models.
Model Type Hugging Face Identifier Use Case
Mistral 7B Pretrained Base Model mistralai/Mistral-7B-v0.1 General language modeling
Chat-Tuned Model mistralai/Mistral-7B-Instruct-v0.1 Dialogue and instruction-following
Math-Tuned Model TIGER-Lab/MAmmoTH2-7B Mathematical reasoning
Code-Tuned Model Nondzu/Mistral-7B-codealpaca-lora Code generation
Table 9: Fine-tuned variants and corresponding Hugging Face identifiers for Mistral 7B models.
Model Type Hugging Face Identifier Use Case
LLaMA2-13B Pretrained Base Model meta-llama/Llama-2-13b-hf General language modeling
Chat-Tuned Model meta-llama/Llama-2-13b-hf Dialogue and instruction-following
Math-Tuned Model TIGER-Lab/MAmmoTH-13B Mathematical reasoning
Code-Tuned Model emre/llama-2-13b-code-chat Code generation
Table 10: Fine-tuned variants and corresponding Hugging Face identifiers for LLaMA-2 13B models.
A.3 Additional Experiment Details
Layer Types on Compression Performance
We further analyze how the choice of compression module impacts performance on LLaMA-2 13B architecture. As shown in Table 11, Chat models exhibit minimal variation in performance across different compression strategies—LLaMA-2 13B favors the attention layers. In contrast, Math and Code models benefit consistently from preserving attention layers, with quantization of these components yielding the smallest performance degradation. This suggests that domain-specific knowledge for complex reasoning tasks is more concentrated in the attention modules, making them less tolerant to aggressive compression.
@@ A.3 Additional Experiment Details
Table 11: Impact of quantizing different module types on LLaMA2-13B. Compression of attention modules yields the best results for Chat, Math and Code models.